OpenStudy (anonymous):

ordered pairs for: x-2y=-14 2x+y=17 HELP ASAP!

jimthompson5910 (jim_thompson5910):

isolate x in the first equation to get x-2y=-14 x=-14+2y

jimthompson5910 (jim_thompson5910):

then plug this into the second equation 2x+y=17 2(-14+2y)+y=17 -28 + 4y + y = 17 5y - 28 = 17 ... ... ... y = ???

OpenStudy (anonymous):

Hmm, I'm still pretty confused.. :/

jimthompson5910 (jim_thompson5910):

do you see how I got what I have so far?

OpenStudy (anonymous):

Yes, I do see. (: So, that's how I'll be finding the ordered pairs?

jimthompson5910 (jim_thompson5910):

the idea from here is to solve 5y - 28 = 17 for y what do you get when you do so?

OpenStudy (anonymous):

9? I'm horrible at math. So, I'm not sure if i done it right but what I did was I subtracted 28 both sides & then I divided it by 5. AGAIN, i'm not sure yet how to solve ordered pairs.............

jimthompson5910 (jim_thompson5910):

y = 9 is correct

jimthompson5910 (jim_thompson5910):

now that you know that y = 9, you can use this to find x x=-14+2y x=-14+2*9 .. plug in y = 9 x = -14 + 18 x = 4

OpenStudy (anonymous):

Can you help me solve x? :)

jimthompson5910 (jim_thompson5910):

so the ordered pair is (4,9)

OpenStudy (anonymous):

Thank you so much!!

jimthompson5910 (jim_thompson5910):

what this means visually is that if you graphed the two lines x-2y=-14 and 2x+y=17, they would intersect at the point (4,9)

jimthompson5910 (jim_thompson5910):

you're welcome